OpenStudy (anonymous):

Find the distance from P to l. Line l contains points (6,5) and (2,3). Point P has coordinates (2,6).

OpenStudy (anonymous):

calculus or geometry?

OpenStudy (anonymous):

geometry

OpenStudy (anonymous):

damn

OpenStudy (anonymous):

ok we need the slope of the line, then the equation for the line perpendicular to the line through the point, and then where they intersect i cannot think of another way to do it with geometry

OpenStudy (anonymous):

slope of the line is easy enough, over 4 up 2 slope is \(\frac{1}{2}\) and so perpendicular line will have slope \(m=-2\)

OpenStudy (anonymous):

equation for the line with slope \(m=-2\) through the point \((2,6)\) is \[y-6=-2(x-2)\] or \[y=-2x+10\]

OpenStudy (anonymous):

equation of the line through the original points is \(y-3=\frac{1}{2}(x-2)\) or \[y=\frac{1}{2}x+2\]

OpenStudy (anonymous):

maybe @tanjung has a snappier way, now i would find where the lines intersect then compute the distance

OpenStudy (anonymous):

ah, thank you very much, that all has helped me.

OpenStudy (anonymous):

find the equation of l through points (6,5) and (2,3), that is x-2y=-4 or x-2y+4=0 then, we calculate the distance P(2,6) to line's l, use formula : d = |(ax1+by1+c)/sqrt(a^2+b^2)| for a,b,c are the contant of line's l and x1,y1 are coordinat's point of P

OpenStudy (anonymous):

thank you very much(: